# Studies of stationary features in jets: 3C 279 quasar I. On-sky scattering and dynamics

### v1

@@ Abstract
Context. A recent study on the dynamics of the quasi-stationary component (QSC) in the jet of BL  Lacertae highlighted its significance in evaluating the physical properties of relativistic transverse waves in the parsec-scale jet. Motivated by this finding, we selected a different type of blazar, the flat-spectrum radio quasar (FSRQ) 3C 279, which hosts a QSC at an angular median distance of 0.35 mas from the radio core, as revealed by 27 years of VLBA monitoring data at 15 GHz.
@@
Methods. We employ an analytical statistical method to estimate the mean intrinsic speed of the QSC, while moving average and refinement methods are used to smooth its trajectory.
Results. Analysis of the QSC position scatter shows that the jet axis change its direction by about $21\degr$ over 27 years and jet mean intrinsic full opening angle is $\approx 0.30\degr\pm 0.03\degr$. The apparent displacement vectors of the QSC exhibit strong asymmetry and anisotropy along the jet direction, indicating pronounced anisotropic displacements of the core along the jet axis. We estimated the mean intrinsic speed of the QSC to be superluminal, $\overline{\beta_{\rm s}}\approx 15$ in units of the speed of light, which, within the framework of the seagull-on-wave model, is interpreted as evidence for a relativistic transverse wave propagating through the QSC. Analysis of the reversing trajectory of the QSC enables the classification and characterization of reversal patterns, which, in turn, allows the determination of key transverse wave parameters such as frequency, amplitude, inclination angle, and magnetic energy of the wave.
Conclusions. Analysing the changes in QSC position and dynamics serves as a powerful diagnostic tool for the jet cone angle and jet direction changes, detecting relativistic transverse waves in the jets of two types of blazars, BL Lacertae and FSRQ 3C 279, and characterizing transverse wave properties.
@@ 1 Introduction
The proximity of the BL Lacertae object (hereafter, BL Lac) and its relatively high radio brightness have motivated a detailed study of the structure and kinematics of its relativistic jet using radio interferometric observations. Cohen et al. (2014, 2015) conducted a comprehensive analysis of the BL Lac jet using VLBA monitoring data at 15 GHz over 17 years. They identified a quasi-stationary component (QSC) located at a projected distance of $\sim 0.26$ mas from the core, along with moving components emanating downstream of the QSC. An examination of the jet ridge line dynamics revealed the presence of transverse waves propagating at relativistic speeds along the jet. Based on the correlated variations in the position angles of the QSC and the transverse wave, they proposed that the transverse waves are excited at the QSC, analogous to shaking the handle of a whip to generate transverse waves. In the ”whip” model, the jet is characterized by a strong helical magnetic field, where the transverse waves are identified as Alfvén waves excited at the QSC, while the moving components represent fast and slow magnetosonic compressions of the magnetic field and/or jet plasma.
@@
Arshakian et al. (2020, 2024) conducted detailed investigations of QSC dynamics in the BL Lac jet, demonstrating that the QSC trajectory exhibits reversal patterns, with apparent speeds frequently exceeding the speed of light. They proposed a model in which the observed reversing motions are projections of the QSC motion across the jet axis, akin to a seagull riding ocean waves. In the ”seagull-on-wave” model, the transverse waves originate upstream of the QSC location, with the QSC moving at subluminal speeds in the host galaxy frame. However, due to projection and relativistic effects, this motion appears to the observer as successive reversing patterns with superluminal speeds. Within this framework, the observed superluminal speeds of the QSC in blazars serve as evidence for the presence of relativistic transverse waves, with the scatter in QSC positions reflecting the amplitude of these waves.
@@
Motivated by these findings in BL Lac, we selected a different type of blazar for further study, the flat-spectrum radio quasar 3C 279, known for its bright, compact radio core at 15 GHz (Lister et al. 2019). The latter is mainly due to the close alignment of the jet axis toward Earth, combined with the high Lorentz factor of the jet beam ($\theta\approx 2\degr$ and $\Gamma_{\rm beam}\approx 20$), which contributes to the observed characteristics of 3C 279 (e.g., Pushkarev et al. 2017). Observations at 15 GHz reveal a QSC located at a median projected distance of 0.35 mas ($2.2$ pc) from the core (Lister et al. 2019). A bright component at about 0.4 mas is visible in the very high angular resolution images of the 3C 279 jet obtained with the space Very Long Baseline Interferometry (VLBI) mission RadioAstron at 22 GHz (Fuentes et al. 2023), believed to correspond to the QSC observed at 15 GHz. At 43 GHz, the QSC appears at shorter separations of 0.25–0.3 mas (Wehrle et al. 2001; Jorstad et al. 2004), which can be due to the variability of this separation between the core and the QSC or to higher angular resolution that resolve the 15 GHz QSC into two features.
@@
We conducted a detailed study of the on-sky distribution of QSC positions and QSC dynamics in the 3C 279 jet using over 25 years of 15 GHz VLBA monitoring data from the MOJAVE program (Monitoring Of Jets in Active Galactic Nuclei with VLBA Experiments, Lister et al. 2009). Observational data and their uncertainties are presented in Section 2. In Section 3, we analyze the on-sky distribution of QSC positions and in Section 4, we examine the QSC kinematics. The characteristics of the QSC trajectory and transverse waves in the jet are estimated in Section 5, followed by the discussion in Section 6.
@@
For the 3C 279 at redshift $z=0.536$ (Marziani et al. 1996), the scale factor is 6.31 pc mas-1, assuming a flat cosmology with $\Omega_{m}=0.27$, $\Omega_{\Lambda}=0.73$, and $H_{0}=71$ km  s-1 Mpc-1 (Komatsu et al. 2009).
@@ 2 Observational data: quasi-stationary component
For the purposes of our study, we used the publicly available VLBA observational data of quasar 3C 279111https://www.cv.nrao.edu/MOJAVE/sourcepages/1253-055.shtml at 15 GHz obtained predominantly within the MOJAVE program (Lister et al. 2021) and its predecessor, the 2-cm VLBA Survey (Kellermann et al. 1998; Zensus et al. 2002), with the rest from the NRAO data archive. The data spans since 1995 July 25 through 2022 March 18, encompassing a total of 167 high-quality observational epochs. The reconstructed brightness distribution maps from these observations typically exhibit a dynamic range of several thousand.
@@
The position and flux density uncertainties of the fitted components were estimated following the approach presented in Schinzel et al. (2012). However, this approach does not provide azimuthal dependence of the positional errors. To account for this, we inferred a median value 0.57 of the ratio of the error across to that along the jet direction following the procedure based on the $\chi^{2}$ statistics as in Arshakian et al. (2020) (see Sec. 3) and Lampton et al. (1976). We tested these theoretical position error estimates by randomly selecting the (u,v)-coverage from 20 epochs and replacing the data in those epochs with known simulated structure based on real models from other randomly selected epochs. The simulated models used a combination of optically thin and thick spheres and varied in number and strength of components. We then conducted a blind modelfit to these simulated data and compared the fitted positions of the innermost component with the known simulated positions. The simulations showed good agreement with the positional errors obtained from the analytical approximations, but required an overall rescaling by 1.37 and additional rescaling to account for their asymmetry.
@@
Figure 1: Scatter of 161 positions of QSC on the sky plane. The sizes of the crosses correspond to QSC position errors in the directions toward and across the core. The radio core is marked by a filled circle at position (0,0), and the median position of the QSC scattering is marked by a red plus sign. The red dashed line is the central axis of the jet, which connects the median positions of the QSC and the core. The green dashed lines represent the linear fit of the cone lines.
@@
The association of the core position with the conical shape of the QSC positions suggests that the latter is due to significant displacements of the core position along the jet axis. The QSC should move almost in a plane normal to the jet axis. In this case, displacements of the core position along the jet axis cause the intrinsic positions of the QSC to scatter downstream and upstream. If the viewing angle of the jet is larger than the intrinsic jet opening angle (i.e. we observe the jet outside its cone generatrix), we should observe an almost conical spread of positions. The size of the scatter along the jet direction reflects the intrinsic motion of the QSC and the motion of the core, while the size of the scatter across the jet through the median center is due to the intrinsic motion of the QSC.
@@
Figure 2: Variation of the QSC position angle with time. The red dashed line is the position angle of the jet central axis, ${\rm PA}_{\rm jet}=-138\degr$, and the blue line is a linear fit to the changes in QSC position angles with time.
@@
Figure 3: The positions of QSC within the six time intervals 1995.57--1999, 1999--2002, 2002--2005.71, 2005.71--2010.9, 2013.24--2020 and 2020--2022.22. The blue plus sign is the median position of the QSC positions within the given time interval. The blue dashed line is the axis of the jet connecting the core and the median position of QSC. The red dashed line is the central axis of the jet, connecting the median position of QSC over the entire observation period 1995.57--2022.22 and the core.
@@ 3 On-sky scatter of QSC positions
The spread of QSC positions on the sky and their position uncertainties are shown in Fig. 1. The median position error across and along the jet is $\sim 5$ $\mu$as and $\sim 12$ $\mu$as, respectively. The small positioning errors are attributed to high compactness and brightness of both the radio core and the quasi-stationary component. The on-sky distribution of the QSC positions has a conical shape with cone lines directed towards the core position at (0,0) (Fig. 1). The size of the scatter along the jet axis is $\sim 0.3$ mas and across the jet at the position of the median center is $\sim 0.15$ mas. To check if the cone lines intersect close to the core and measure the cone angle, we excluded three positions at $\rm{R.A.}>-0.12$ mas, which are positions of the last three observational epochs (see Fig. 2) and stand out from the main scatter of positions. We used binning of positions in six non-overlapping bands transverse to the jet axis; for each band, we selected positions in the 85% quintile on either side of the jet axis and performed a linear fit for the selected positions on either side. The intersection of the conical lines (the apex of the cone) is located very close to the core at a distance of 0.02 mas. According to our estimates, the cone angle is about $29\degr$. A change in the quintile from 75% to 90% results in a change in the angle of the cone that ranges from $26\degr$ to $32\degr$, respectively, and the distance between the core and apex of the cone within 0.05 mas. We assume that the conical shape of the position distribution is associated with the VLBI core and take the apparent full-cone angle to be $\phi_{\rm app}\approx 30\degr$. The latter corresponds approximately to the ±$\pm 1.5$ standard deviation of the QSC position angles, $\sigma_{\rm PA_{\rm s}}$. (Further, all designations with subscripts $s$ or $c$ indicate the intrinsic characteristics of the QSC and the core, respectively). In the following, we will use the value $2\times(1.5\sigma_{\rm PA_{\rm s}})=3\sigma_{\rm PA_{\rm s}}$ as a proxy for the apparent full cone angle.
@@
The direction of the jet axis, as defined by PAs, exhibits a general trend of moving southward. Initially, the jet axis moves clockwise until 2005.71, then reverses direction during the period 2005.71 to 2010.9, before resuming its clockwise motion toward the south. This behavior may result from the superposition of two precessional motions: a long-term precession that drives the overall southward shift and a shorter-term precession (nutation) with a characteristic timescale of approximately five years (as observed between 2005.71 and 2010.9). Over a span of 27 years, the jet axis has changed its orientation on the sky by $21\degr$. If this variation is due to long-term jet precession, the corresponding precession angle and period must be at least $21\degr$ and $\gtrsim 40$ years, respectively. Extended VLBA monitoring of 3C 279 at 15 GHz is required to test the jet precession hypothesis. The change in PA of the inner jet towards the South direction was also revealed from the RadioAstron observations in March 2014 (Fuentes et al. 2023) and February 2018 (Toscano et al., subm.).
The average apparent cone angle calculated over six time intervals is $\overline{\phi}_{\rm app}=15.7\degr\pm 1.6\degr$. This estimate closely matches twice the residual standard error of the linear fit shown in Fig. 2, with $15.7\approx 2\times 7.1=14.2$. Significant deviations of $\phi_{\rm app}$ from the mean value occur during the period 1999–2005.71. Our estimate of the mean apparent cone angle is consistent with other measurements. An apparent cone angle of $\phi_{\rm app}\approx 12\degr$ was measured by modeling the size of the core and the moving C4 component at 22 GHz (Wehrle et al. 2001), using VLBA monitoring data from 1992 to 1998. This time interval partially overlaps with the period 1995.57–1999, during which we obtained $\phi_{\rm app}=15\degr$. The apparent cone angles of $\phi_{\rm app}=14\degr$ and $\phi_{\rm app}=11.6\degr\pm 0.8\degr$ were also estimated using a single-epoch 15 GHz map (Pushkarev et al. 2009) and multi-epoch 15 GHz VLBA observations, respectively, with the latter based on deconvolution of the transverse jet width along the jet ridge line at distances $>0.5$ mas from the core (Pushkarev et al. 2017). We argue that all these estimates are valid and that the differences between them are likely due to the use of data from different observational epochs and recognize that the apparent opening angle of the QSC cone tracks well with the apparent opening angle of the jet, as determined by the moving components and the transverse width of the jet.
@@
In the seagull-on-wave model, the QSC moves nearly across the jet axis due to the passage of a relativistic, swirled transverse waves (Arshakian et al. 2024), and the extent of the QSC scatter reflects the maximum amplitude of this wave. The transverse size of the opening cone at the median position, $d_{\rm t}=d_{\rm j}\,\phi_{\rm app}$, where $d_{\rm j}$ is the distance along the jet axis to the median position of QSC, can thus serve as an indicator of the maximum amplitude of the transverse wave in the jet, $A_{\rm w,max}$. For the six specified time intervals, we estimated $d_{\rm t}\approx A_{\rm w,max}$ values of 0.071, 0.069, 0.129, 0.114, 0.107, and 0.082 mas (Fig. 2). The maximum amplitude of the transverse wave varies almost twice from 0.07 mas to 0.13 mas, with a mean maximum amplitude of $\overline{A}_{\rm w,max}=0.10\pm 0.01$ mas ($0.60\pm 0.06$ pc).

### v2

@@ Abstract
Context. A recent study on the dynamics of the quasi-stationary component (QSC) in the jet of BL Lacertae highlighted its significance in evaluating the physical properties of relativistic transverse waves in the parsec-scale jet. Motivated by this finding, we selected a different type of blazar, the flat-spectrum radio quasar (FSRQ) 3C 279, which hosts a QSC at an angular median distance of 0.35 mas from the radio core, as has been revealed by 27 years of VLBA monitoring data at 15 GHz.
@@
Methods. We employed an analytical statistical method to estimate the mean intrinsic speed of the QSC, while moving average and refinement methods were used to smooth its trajectory.
Results. Analysis of the QSC position scatter shows that the jet axis changes its direction by about $21\degr$ over 27 years and jet mean intrinsic full opening angle is $\approx 0.30\degr\pm 0.03\degr$. The apparent displacement vectors of the QSC exhibit strong asymmetry and anisotropy in the direction of the jet, indicating pronounced anisotropic displacements of the core along the jet axis. We estimated the mean intrinsic speed of the QSC to be superluminal, $\overline{\beta_{\rm s}}\approx 10$ in units of the speed of light, which, within the framework of the seagull-on-wave model, is interpreted as evidence of a relativistic transverse wave propagating through the QSC. Analysis of the reversing trajectory of the QSC enables the classification and characterization of reversal patterns, which, in turn, allows us to determine key transverse wave parameters such as the frequency, amplitude, inclination angle, and magnetic energy of the wave.
Conclusions. Analysing the variations in the QSC position and dynamics provides a powerful diagnostic for determining changes in the jet cone angle and direction, for detecting relativistic transverse waves in the jets of two types of blazars, BL Lacertae and the FSRQ 3C 279, and for characterising the properties of these transverse waves.
@@ 1 Introduction
The proximity of the BL Lacertae object (hereafter, BL Lac) and its relatively high radio brightness have motivated a detailed study of the structure and kinematics of its relativistic jet using radio interferometric observations. cohen14; cohen15 conducted a comprehensive analysis of the BL Lac jet using VLBA monitoring data at 15 GHz over 17 years. They identified a quasi-stationary component (QSC) located at a projected distance of $\sim 0.26$ mas from the core, along with moving components emanating downstream of the QSC. An examination of the jet ridge line dynamics revealed the presence of transverse waves propagating at relativistic speeds along the jet. Based on the correlated variations in the position angles of the QSC and the transverse wave, they proposed that the transverse waves are excited at the QSC, analogous to shaking the handle of a whip to generate transverse waves. In the ‘whip’ model, the jet is characterized by a strong helical magnetic field, where the transverse waves are identified as Alfvén waves excited at the QSC, while the moving components represent fast and slow magnetosonic compressions of the magnetic field and/or jet plasma.
@@
arshakian20; arshakian24 conducted detailed investigations of QSC dynamics in the BL Lac jet, demonstrating that the QSC trajectory exhibits reversal patterns, with apparent speeds frequently exceeding the speed of light. They proposed a model in which the observed reversing motions are projections of the QSC motion across the jet axis, akin to a seagull riding ocean waves. In the ‘seagull-on-wave’ model, the transverse waves originate upstream of the QSC location, with the QSC moving at subluminal speeds in the host galaxy frame. However, due to projection and relativistic effects, this motion appears to the observer as successive reversing patterns with superluminal speeds. Within this framework, the observed superluminal speeds of the QSC in blazars serve as evidence of the presence of relativistic transverse waves, with the scatter in QSC positions reflecting the amplitude of these waves.
@@
Motivated by these findings in BL Lac, we selected a different type of blazar for further study, the flat-spectrum radio quasar (FSRQ) 3C 279, known for its bright, compact radio core at 15 GHz (lister19). The latter is mainly due to the close alignment of the jet axis towards Earth, combined with the high Lorentz factor of the jet beam ($\theta\approx 2\degr$ and $\Gamma_{\rm beam}\approx 20$), which contributes to the observed characteristics of 3C 279 (e.g. pushkarev17). Observations at 15 GHz reveal a QSC located at a median projected distance of 0.35 mas ($2.2$ pc) from the core (lister19). A bright component at about 0.4 mas is visible in the very high-angular-resolution images of the 3C 279 jet obtained with the space Very Long Baseline Interferometry (VLBI) mission RadioAstron at 22 GHz (fuentes23), believed to correspond to the QSC observed at 15 GHz. At 43 GHz, the QSC appears at shorter separations of 0.25–0.3 mas (wehrle01; jorstad04), which could be due to the variability of this separation between the core and the QSC or to higher angular resolution that resolve the 15 GHz QSC into two features.
@@
We conducted a detailed study of the on-sky distribution of QSC positions and QSC dynamics in the 3C 279 jet using over 25 years of 15 GHz VLBA monitoring data from the MOJAVE programme (Monitoring Of Jets in Active Galactic Nuclei with VLBA Experiments, Lister et al. 2009). Observational data and their uncertainties are presented in Section 2. In Section 3, we analyse the on-sky distribution of QSC positions, and in Section 4, we examine the QSC kinematics. The characteristics of the QSC trajectory and transverse waves in the jet are estimated in Section 5, followed by the discussion in Section 6.
@@
We assumed a flat cosmology with $\Omega_{m}=0.27$, $\Omega_{\Lambda}=0.73$, and $H_{0}=71$ km s-1 Mpc-1 (komatsu09). For the 3C 279 at redshift $z=0.536$ (marziani96), the scale factor is 6.31 pc mas-1.
@@ 2 Observational data: Quasi-stationary component
For the purposes of our study, we used the publicly available VLBA observational data of quasar 3C 279111https://www.cv.nrao.edu/MOJAVE/sourcepages/1253-055.shtml at 15 GHz obtained predominantly within the MOJAVE programme (lister21) and its predecessor, the 2-cm VLBA Survey (kellermann98; Zensus02), with the rest from the NRAO data archive. The data spans from July 25 1995 through March 18 2022, encompassing a total of 167 high-quality observational epochs. The reconstructed brightness distribution maps from these observations typically exhibit a dynamic range of several thousand.
@@
The position and flux density uncertainties of the fitted components were estimated following the approach presented in schinzel12. However, this approach does not provide azimuthal dependence of the positional errors. To account for this, we inferred a median value of 0.57 for the ratio of the error across of the jet to that in the direction of the jet, following the procedure based on the $\chi^{2}$ statistics described in arshakian20 (see Sec. 3) and lampton76. We tested these theoretical position error estimates by randomly selecting the (u,v) coverage from 20 epochs and replacing the data in those epochs with known simulated structure based on real models from other randomly selected epochs. The simulated models used a combination of optically thin and thick spheres and varied in terms of the number and strength of components. We then conducted a blind modelfit to these simulated data and compared the fitted positions of the innermost component with the known simulated positions. The simulations showed good agreement with the positional errors obtained from the analytical approximations, but required an overall rescaling by 1.37 and additional rescaling to account for their asymmetry.
@@
Figure 1: Scatter of 161 positions of QSC on the sky plane. The sizes of the crosses correspond to QSC position errors in the directions toward of and across the core. The radio core is marked by a filled circle at position (0,0), and the median position of the QSC scattering is marked by a red plus sign. The dashed red line is the central axis of the jet, which connects the median positions of the QSC and the core. The dashed green lines represent the linear fit of the cone lines.
@@
The association of the core position with the conical shape of the QSC positions suggests that the latter is due to significant displacements of the core position along the jet axis. The QSC should move almost in a plane normal to the jet axis. In this case, displacements of the core position along the jet axis cause the intrinsic positions of the QSC to scatter downstream and upstream. If the viewing angle of the jet is larger than the intrinsic jet opening angle (i.e. we observe the jet outside its cone generatrix), we should observe an almost conical spread of positions. The size of the scatter in the direction of the jet reflects the intrinsic motion of the QSC and the motion of the core, while the size of the scatter across the jet through the median centre is due to the intrinsic motion of the QSC.
@@
Figure 2: Variation in QSC position angle with time. The dashed red line is the position angle of the jet central axis, ${\rm PA}_{\rm jet}=-138\degr$, and the blue line is a linear fit to the changes in QSC position angles with time.
@@
Figure 3: Positions of the QSC within the six time intervals, 1995.57-1999, 1999-2002, 2002-2005.71, 2005.71-2010.9, 2013.24-2020 and 2020-2022.22. The blue plus sign is the median position of the QSC positions within the given time interval. The dashed blue line is the axis of the jet connecting the core and the median position of QSC. The dashed red line is the central axis of the jet, connecting the median position of QSC over the entire observation period 1995.57-2022.22 and the core. The sizes of the crosses correspond to QSC position errors in the directions toward of and across the core.
@@ 3 On-sky scatter of QSC positions
[…]
The direction of the jet axis, as defined by PAs, exhibits a general trend of moving southwards. Initially, the jet axis moves clockwise until 2005.71, then reverses direction during the period 2005.71 to 2010.9, before resuming its clockwise motion towards the south. This behaviour may result from the superposition of two precessional motions: a long-term precession that drives the overall southward shift and a shorter-term precession (nutation) with a characteristic timescale of approximately five years (as observed between 2005.71 and 2010.9). Over a span of 27 years, the jet axis has changed its orientation on the sky by $21\degr$. If this variation is due to long-term jet precession, the corresponding precession angle and period must be at least $21\degr$ and $\gtrsim 40$ years, respectively. The position angles at the last three epochs suggest that the jet axis continues the long-term trend of its southward motion (Fig. 2). Extended VLBA monitoring of 3C 279 at 15 GHz is required to test the jet precession hypothesis. The change in the PA of the inner jet towards a southerly direction was also revealed from the RadioAstron observations in March 2014 (fuentes23) and February 2018 (toscano25).
The average apparent cone angle calculated over six time intervals is $\overline{\phi}_{\rm app}=15.7\degr\pm 1.6\degr$. This estimate closely matches twice the residual standard error of the linear fit shown in Fig. 2, with $15.7\approx 2\times 7.1=14.2$. Significant deviations of $\phi_{\rm app}$ from the mean value occur during the period 1999–2005.71. Our estimate of the mean apparent cone angle is consistent with other measurements. An apparent cone angle of $\phi_{\rm app}\approx 12\degr$ was measured by modelling the size of the core and the moving C4 component at 22 GHz (wehrle01), using VLBA monitoring data from 1992 to 1998. This time interval partially overlaps with the period 1995.57–1999, during which we obtained $\phi_{\rm app}=15\degr$. The apparent cone angles of $\phi_{\rm app}=14\degr$ and $\phi_{\rm app}=11.6\degr\pm 0.8\degr$ were also estimated using a single-epoch 15 GHz map (pushkarev09) and multi-epoch 15 GHz VLBA observations, respectively, with the latter based on deconvolution of the transverse jet width along the jet ridge line at distances >0.5>0.5 mas from the core (pushkarev17). We argue that all these estimates are valid and that the differences between them are likely due to the use of data from different observational epochs and recognize that the apparent opening angle of the QSC cone tracks well with the apparent opening angle of the jet, as determined by the moving components and the transverse width of the jet.
@@
In the seagull-on-wave model, the QSC moves nearly across the jet axis due to the passage of a relativistic, swirled transverse waves (arshakian24), and the extent of the QSC scatter reflects the maximum amplitude of this wave. The transverse size of the opening cone at the median position, $d_{\rm t}=d_{\rm j}\,\phi_{\rm app}$, where $d_{\rm j}$ is the distance along the jet axis to the median position of QSC, can thus serve as an indicator of the maximum amplitude of the transverse wave in the jet, $A_{\rm w,max}$. For the six specified time intervals, we estimated $d_{\rm t}\approx A_{\rm w,max}$ values of 0.071, 0.069, 0.129, 0.114, 0.107, and 0.082 mas (Fig. 2). The maximum amplitude of the transverse wave varies almost twice from 0.07 mas to 0.13 mas, with a mean maximum amplitude of $\overline{A}_{\rm w,max}=0.10\pm 0.01$ mas ($0.60\pm 0.06$ pc).